# Accelerating the Exploration of Thermal Materials via a Synergistic Strategy of DFT and Interpretable Deep Learning

### v9

@@ II.3 Screening potential thermal insulators/conductors
According to our MD calculations, the actual $\kappa_{\rm L}$ values of the candidate materials fall within distinct ranges ($\sim$30, $\sim$80, $\sim$150 W/m/K). As shown in Figure 8(e), samples with $\kappa_{\rm L}>100$ W/m/K account for only 0.47% of the entire training dataset. In such cases of data imbalance, model predictions for regions with scarce samples typically exhibit considerable uncertainty [76, 77].
KRbAu2 has the space group symbol $\textit{Fm}\overline{3}\textit{m}$, belonging to the cubic crystal system. For such isotropic thermal insulator candidates, it is generally more efficient and accurate to calculate $\kappa_{\rm L}$ using DFT [78, 79]. As shown in Figure 8(d), the DFT-calculated $\kappa_{\rm L}$ for this structure at 300K is 0.201 W/m/K, while the KAN-predicted value is 0.2169, correctly identifying it as a thermal insulator. In summary, the methods of $k$-means clustering and t-SNE visualization are also sufficient for qualitatively assessing the $\kappa_{\rm L}$ of materials, enabling high-throughput screening in combination with the pretrained model. Furthermore, the $\kappa_{\rm L}$ values predicted by KAN exhibit excellent agreement with both MD and DFT validation results. This demonstrates that the pretrained KAN model achieves not only reliable accuracy but also satisfactory extrapolation capability.
II.4 Tracing high/low LTC from the perspective of theoretical calculation
Among the validated candidates, considering the isotropy of Be2C and KRbAu2, we take these two as examples to explore the mechanisms influencing high and low $\kappa_{\rm L}$ based on DFT. Their crystal structures are shown in Figures 9(a) and (b), both belonging to the cubic system but exhibiting distinctly different phonon transport behaviors. Be2C adopts a [BeC4] tetrahedral configuration, where each C atom is coordinated with four Be atoms, with a bond length of approximately 1.64 ${\rm\AA}$. These [BeC4] tetrahedral units share edges, forming a continuous three-dimensional covalent network that imparts high rigidity and thermal stability to the crystal. In contrast, KRbAu2 consists of K, Rb, and Au, forming a metal-ion mixed-bond network. The Au atoms exist as [Au2] dimers, where each Au atom is connected via metallic bonds to form a framework structure. K and Rb act as cations filling the interstitial spaces, providing electrostatic stabilization. Due to the large atomic radii of K and Rb, the introduced steric hindrance effect increases lattice flexibility [80].
Figure 9: Based on DFT-based phonon thermal transport analysis. Crystal structures and the projected 2D electron localization function (ELF) diagram of (a) CrB4 and (b) KRbAu2. (c) is the -pCOHP & -IpCOHP for Be-C, K-Au and Rb-Au bonds. In (d) and (e), the left panels depict the phonon dispersion curves of CrB4 and KRbAu2, while the right panels present the corresponding atom-projected PDOS.
@@ II.4 Tracing high/low LTC from the perspective of theoretical calculation
Figure 10: Modal quantities of CrB4 and KRbAu2. (a-c) illustrate the frequency dependence of the average thermal conductivity $\bar{\kappa}_{\lambda}$, group velocity $|\mathbf{v}_{\lambda}|$, and phonon lifetime $\tau_{\lambda}$ for CrB4, respectively; (d-f) present the counterparts for KRbAu2. The scatters are colored according to the band index, with two distinct colormaps used to differentiate the two materials.
@@
To further investigate the impact of chemical bonding on the $\kappa_{\rm L}$ of Be2C and KRbAu2, we employ the Negative Projected Crystal Orbital Hamilton Population (-pCOHP) to visualize their local bonding characteristics. In -pCOHP, positive values indicate bonding states, negative values represent antibonding states, and a value of zero signifies non-bonding interactions. As shown in Figure 9(c), the Fermi level $E_{\rm F}$ lies at 0 eV. For the Be-C bond, bonding states dominate below $E_{\rm F}$ down to -8.7 eV. Strong bonding implies greater lattice rigidity and a steeper potential energy surface [87], leading to higher optical branch vibration frequencies, which suppress phonon scattering and enhance thermal conductivity. This behavior is similar to classical high-$\kappa_{\rm L}$ materials such as diamond and SiC. In KRbAu2, the bonding characteristics differ between the two types of bonds: K-Au is bonding, while Rb-Au is antibonding. The antibonding nature of the Rb-Au bond softens the lattice, enhances phonon scattering, and reduces $\kappa_{\rm L}$. Figure 8(d) confirms the ultra-low $\kappa_{\rm L}$ of KRbAu2, suggesting that its phonon transport properties are primarily dictated by the Rb-Au bond. The -IpCOHP analysis further evaluates global bond strength. Despite K-Au being a bonding interaction, its bond strength remains weak, and the presence of an antibonding Rb-Au bond exacerbates phonon anharmonicity. Weak chemical bonds are often associated with strong phonon anharmonicity [88, 89], which is a key factor underlying the ultra-low $\kappa_{\rm L}$ of KRbAu2.
@@
Figures 9(d) and (e) provide insights into the characteristics of low-energy optical branches and acoustic modes, supporting the previous hypotheses. In the phonon dispersion spectrum of Be2C, all phonon modes exhibit real frequencies, indicating dynamical stability at zero temperature with no tendency for structural collapse. The acoustic branches in the low-frequency region (0 $\sim$ 10 THz) show relatively high group velocities, suggesting potentially high phonon thermal conductivity. In contrast, the optical branches are primarily distributed in the high-frequency range (15 $\sim$ 30 THz) and exhibit a certain degree of flatness, indicating lower group velocities and minimal contribution to heat transport. The Projected Density of State (PDOS) analysis on the right side of Figure 9(d) further reveals that heat-carrying phonons are predominantly below 15 THz, mainly contributed by the heavier Be atoms, while the high-frequency optical branches involve both C and Be atoms. For KRbAu2, the phonon dispersion curve shows that all modes remain within the real frequency range, with no imaginary branches, confirming its dynamical stability at zero temperature. The acoustic branches are concentrated in the low-frequency region (0 $\sim$ 1 THz) and exhibit flat dispersions, resulting in low group velocities. The optical branches are distributed between 1 and 3.5 THz, with some modes displaying significant dispersion. According to the PDOS, the dominant vibrating atoms shift sequentially from Au to Rb to K as frequency increases, consistent with the mass distribution of these elements.
@@
Figure 10(b) and (e) show the frequency dependence of the group velocity $\mathbf{v}_{g}$ at the center of the Brillouin zone for Be2C and KRbAu2, respectively. The group velocity of Be2C exceeds 10 km/s, while that of KRbAu2 does not exceed 1.1 km/s. This significant difference in group velocity can be attributed to differences in bonding and strength, which is consistent with the inference made from the steepness of their phonon spectra. Figure 10(c) and (f) describe the differences in phonon lifetimes between the two materials. In the low-frequency range below 10 THz, Be2C exhibits longer phonon lifetimes, with the longest exceeding 10 ps and the shortest being above 1 ps. In contrast, in KRbAu2, the phonon lifetime decreases sharply with increasing frequency, with $\tau_{\lambda}$ approaching 0.1 ps at $\upsilon_{\lambda}=2{\rm THz}$. It is noteworthy that for certain specific low-frequency modes, the phonon lifetime $\tau_{\lambda}$ of KRbAu2 exceeds that of Be2C, which may be due to special scattering suppression effects, e.g., weaker three-phonon scattering [90]. However, overall, Be2C has a longer and more uniform phonon lifetime, which is favorable for long-range heat transport, while the shorter phonon lifetime of KRbAu2 suppresses $\kappa_{\rm L}$.
III DISCUSSION
@@
Our work employed KAN to model $\log(\kappa_{\rm L})$, demonstrating that its performance in terms of both accuracy and robustness is fully comparable to that of black-box models. For the interpretability analysis of white-box models, we considered feature interaction effects and dependencies. Sensitivity analysis results based on the KAN and SISSO models revealed that KAN can accurately extract key features of $\kappa_{\rm L}$, highlighting its superior interpretability in this context. We combined KAN with CGCNN to construct a two-stage prediction framework, where CGCNN predicts key features, and the KAN pretraining model maps these features to $\log(\kappa_{\rm L})$; for physical properties like $C_{\text{v}}$, which CGCNN struggles to predict accurately, we used DFT calculations to complement the framework. In the case of unlabeled new samples, we applied $k$-means clustering to identify 10 potential thermal conductors/insulators. Among them, the $\kappa_{\rm L}$ values of CN2, CrB4, Be2C, and KRbAu2 were verified by DFT and MD, yielding results highly consistent with deep learning predictions, which sufficiently demonstrate KAN’s excellent extrapolation capability. We visualized the phonon spectra, PDOS, -pCOHP, and other information of Be2C and KRbAu2 using DFT, analyzing the phonon transport mechanisms of both materials from the perspectives of chemical bonding, atomic mass, and structural stability, providing theoretical insights as reference.
IV METHODS
IV.1 Kolmogorov-Arnold Networks (KANs)
where ${{\phi}_{q,p}}:\left[0,1\right]\to\mathbb{R}$ as well as ${{\Phi}_{q}}:\mathbb{R}\to\mathbb{R}$. In this context, $p$ denotes the number of top operators, $q$ denotes the number of bottom operators, and $n$ denotes the number of nodes in the bottom network (which corresponds to the dimensionality of the input features in the input layer), $\mathbf{x}=\left({{x}_{1}},...,{{x}_{n}}\right)$ represents feature vector.
In the original K-A representation theorem, the number of nonlinear layers is limited to 2, and the number of hidden layer nodes is set to $12n+$ 1, which fixes the network structure to $\left[n,2n+1,1\right]$. However, Liu et al. [30] are not constrained by these limitations. In KANs, both the number of layers and the width of the network are arbitrary, which enhances the feasibility of the K-A representation theorem for ML applications. Therefore, for a KAN with ${{n}_{0}}$-D inputs, L layers, and ${{n}_{\text{L}}}=1$-D output, a more precise definition is:
the expression can be simplified to:
where B-spline function is defined as:
$B_{i,0}(x)$ =$=\begin{cases}1,&\text{if }t_{i}\leq x<t_{i+1}\\ 0,&\text{otherwise}\end{cases}$
$B_{i,k}(x)$ $=\frac{x-t_{i}}{t_{i+k}-t_{i}}B_{i,k-1}(x)+\frac{t_{i+k+1}-x}{t_{i+k+1}-t_{i+1}}B_{i+1,k-1}(x)$ (11)
IV.2 Sure Independence Screening & Sparsifying Operator (SISSO)
${{\hat{H}}^{\left(\text{m}\right)}}\equiv\left\{I,+,-,\times,\div,\exp,\log,\text{abs},\sqrt{{}}{{,}^{-1}}{{,}^{-2}}{{,}^{-3}}\right\}\left[{{\phi}_{1}},{{\phi}_{2}}\right]$ (12)
${{\Phi}_{i}}=\bigcup\limits_{k}{{{{\hat{h}}}_{k}}^{\left(\text{m}\right)}\left[{{{\hat{\phi}}}_{i}},{{{\hat{\phi}}}_{j}}\right]},\forall{{\hat{h}}_{k}}^{\left(\text{m}\right)}\in{{\hat{H}}^{\left(\text{m}\right)}}\text{ and }\forall{{\phi}_{i}},{{\phi}_{j}}\in{{\Phi}_{i-1}}$ (13)
$\mathbf{\hat{P}}={{\mathbf{d}}_{n}}{{\mathbf{c}}_{n}}={{c}_{0}}+{{c}_{1}}{{d}_{1}}+\cdots+{{c}_{n}}{{d}_{n}}$ (14)
IV.3 Sobol Indices
${{S}_{i}}=\frac{\text{Var}\left[{{\text{E}}_{\sim i}}\left[Y|{{X}_{i}}\right]\right]}{\text{Var}\left[Y\right]}$ (15)
$S_{i}^{\text{T}}=\frac{{{\text{E}}_{\sim i}}\left[\text{Var}\left[Y|{{\mathbf{X}}_{\sim i}}\right]\right]}{\text{Var}\left[Y\right]}$ (16)
IV.4 Kucherenko Indices
Unlike metrics such as Sobol, LIME, and SHAP, the Kucherenko indices accounts for the dependencies among input features. As an extension of the Sobol sensitivity indices, the Kucherenko indices is specifically designed to quantify the sensitivity of model outputs to input variables while considering these interdependencies [62, 65, 94]. The Kucherenko indices essentially involves computing Sobol indices after first employing Copulas [95] (typically Gaussian Copulas) to separate each feature’s marginal distribution from the dependency structure [96], as detailed in Table S2. This approach approximates the correlated variables as independent before calculating the Sobol indices. The Gaussian copula is constructed based on the standard normal distribution, and its mathematical expression is:
$C\left({{u_{1}},{u_{2}},\ldots,{u_{d}};\Sigma}\right)={\Phi_{\Sigma}}\left[{{\Phi^{-1}}\left({{u_{1}}}\right),{\Phi^{-1}}\left({{u_{2}}}\right)\ldots,{\Phi^{-1}}\left({{u_{d}}}\right)}\right]$ (17)
IV.5 Dataset construction & feature preprocessing
$x^{\prime}=\frac{{x-{x_{\min}}}}{{{x_{\max}}-{x_{\min}}}}$ (18)
IV.6 Implementation of DFT & GPUMD methods
DFT and AIMD calculations are performed using VASP [104, 105]. The Projector-Augmented Wave (PAW) method [106] and the Perdew-Burke-Ernzerhof (PBE) functional [107] within the Generalized Gradient Approximation (GGA) are employed for electron exchange-correlation. The cutoff energy is set to 500 eV, and the electronic convergence threshold is $10^{-8}$ eV.
For the calculation of $\kappa_{\rm L}$, both the phonon BTE method and MD simulations are used. For the stable structures, $\kappa_{\rm L}$ are computed using phono3py [108, 109, 110]. While $\kappa_{\rm L}$ of unstable structures are determined through GPUMD [111] AIMD-NVT simulations are carried out in the canonical ensemble (NVT) with a Nosé–Hoover thermostat at 300 K for a duration of 5 ns. Uniformly spaced sampling is performed, with 250 frames extracted to train the NEP model. The HNEMD method is employed in GPUMD simulations. A Fe value of $1\times 10^{-4}\,{\rm\AA}^{-1}$ has been tested and found to be optimal for this study. Specific heat and phonon dispersions are calculated using Phonopy [112]. The pCOHP analysis is performed using the LOBSTER code [113]. Structural visualizations and ELF analysis are conducted using the VESTA software [114].
IV.7 Data Availability
@@
The original LTC data and the selected features can be downloaded from https://aflowlib.org/ or accessed via the GitHub homepage at https://github.com/FlorianTseng/LTC-modeling.
IV.8 Code Availability
@@
IV.9 Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgements.
We would like to acknowledge the National Key Research and Development Program of China (Grant No. 2023YFB4603800), the financial support from the project of the National Natural Science Foundation of China (Grants No.12404062, No.12474093, No.12302220, No.12374091), Translational Medicine and Interdisciplinary Research Joint Fund of Zhongnan Hospital of Wuhan University (Grant NO.ZNJC202235), and the Fundamental Research Funds for the Central Universities (Grant No.2042023kf0109). The numerical calculations in this work have been done on the supercomputing system in the Supercomputing Center of Wuhan University.

### v10

@@ III.3 Screening potential thermal insulators/conductors
According to our MD calculations, the actual $\kappa_{\rm L}$ values of the candidate materials fall within distinct ranges ($\sim 30$, $\sim 80$, $\sim 150$ $\rm{W\cdot m^{-1}\cdot K^{-1}}$). As shown in Fig. 10(e), samples with $\kappa_{\rm L}>100$ $\rm{W\cdot m^{-1}\cdot K^{-1}}$ account for only 0.47% of the entire training dataset. In such cases of data imbalance, model predictions for regions with scarce samples typically exhibit considerable uncertainty [107, 108].
Figure 11: Based on DFT-based phonon thermal transport analysis. Crystal structures and the projected 2D electron localization function (ELF) diagram of (a) Be2C and (b) KRbAu2. (c) is the -pCOHP & -IpCOHP for Be-C, K-Au and Rb-Au bonds. In (d) and (e), the left panels depict the phonon dispersion curves of Be2C and KRbAu2, while the right panels present the corresponding atom-projected PDOS.
KRbAu2 has the space group symbol $\textit{Fm}\overline{3}\textit{m}$, belonging to the cubic crystal system. For such isotropic thermal insulator candidates, it is generally more efficient and accurate to calculate $\kappa_{\rm L}$ using DFT [109, 110]. As shown in Fig. 10(d), the DFT-calculated $\kappa_{\rm L}$ for this structure at 300K is 0.201 $\rm{W\cdot m^{-1}\cdot K^{-1}}$, while the KAN-predicted value is 0.2169, correctly identifying it as a thermal insulator. In summary, the methods of $k$-means clustering and $t$-SNE visualization are also sufficient for qualitatively assessing the $\kappa_{\rm L}$ of materials, enabling high-throughput screening in combination with the pretrained model. Furthermore, the $\kappa_{\rm L}$ values predicted by KAN exhibit excellent agreement with both MD and DFT validation results. This demonstrates that the pretrained KAN model achieves not only reliable accuracy but also satisfactory extrapolation capability.
III.4 Tracing high/low LTC from the perspective of theoretical calculation
[…]
Figure 12: Modal quantities of Be2C and KRbAu2. (a-c) illustrate the frequency dependence of the average thermal conductivity $\bar{\kappa}_{\lambda}$, group velocity $|\mathbf{v}_{\lambda}|$, and phonon lifetime $\tau_{\lambda}$ for Be2C, respectively; (d-f) present the counterparts for KRbAu2. The scatters are colored according to the band index, with two distinct colormaps used to differentiate the two materials.
@@ III.4 Tracing high/low LTC from the perspective of theoretical calculation
To further investigate the impact of chemical bonding on the $\kappa_{\rm L}$ of Be2C and KRbAu2, we employ the Negative Projected Crystal Orbital Hamilton Population (-pCOHP) to visualize their local bonding characteristics. In -pCOHP, positive values indicate bonding states, negative values represent antibonding states, and a value of zero signifies non-bonding interactions. As shown in Fig. 11(c), the Fermi level $E_{\rm F}$ lies at 0 eV. For the Be-C bond, bonding states dominate below $E_{\rm F}$ down to -8.7 eV. Strong bonding implies greater lattice rigidity and a steeper potential energy surface [118], leading to higher optical branch vibration frequencies, which suppress phonon scattering and enhance thermal conductivity. This behavior is similar to classical high-$\kappa_{\rm L}$ materials such as diamond and SiC. In KRbAu2, the bonding characteristics differ between the two types of bonds: K-Au is bonding, while Rb-Au is antibonding. The antibonding nature of the Rb-Au bond softens the lattice, enhances phonon scattering, and reduces $\kappa_{\rm L}$. Fig. 10(d) confirms the ultra-low $\kappa_{\rm L}$ of KRbAu2, suggesting that its phonon transport properties are primarily dictated by the Rb-Au bond. The -IpCOHP analysis further evaluates global bond strength. Despite K-Au being a bonding interaction, its bond strength remains weak, and the presence of an antibonding Rb-Au bond exacerbates phonon anharmonicity. Weak chemical bonds are often associated with strong phonon anharmonicity [119, 120], which is a key factor underlying the ultra-low $\kappa_{\rm L}$ of KRbAu2.
@@
Figures 11(d) and (e) provide insights into the characteristics of low-energy optical branches and acoustic modes, supporting the previous hypotheses. In the phonon dispersion spectrum of Be2C, all phonon modes exhibit real frequencies, indicating dynamical stability at zero temperature with no tendency for structural collapse. The acoustic branches in the low-frequency region ($0\sim 10$ THz) show relatively high group velocities, suggesting potentially high phonon thermal conductivity. In contrast, the optical branches are primarily distributed in the high-frequency range (15 $\sim$ 30 THz) and exhibit a certain degree of flatness, indicating lower group velocities and minimal contribution to heat transport. The Projected Density of State (PDOS) analysis on the right side of Fig. 11(d) further reveals that heat-carrying phonons are predominantly below 15 THz, mainly contributed by the heavier Be atoms, while the high-frequency optical branches involve both C and Be atoms. For KRbAu2, the phonon dispersion curve shows that all modes remain within the real frequency range, with no imaginary branches, confirming its dynamical stability at zero temperature. The acoustic branches are concentrated in the low-frequency region (0 $\sim$ 1 THz) and exhibit flat dispersions, resulting in low group velocities. The optical branches are distributed between 1 and 3.5 THz, with some modes displaying significant dispersion. According to the PDOS, the dominant vibrating atoms shift sequentially from Au to Rb to K as frequency increases, consistent with the mass distribution of these elements.
@@
Fig. 12(b) and (e) show the frequency dependence of the group velocity $\mathbf{v}_{g}$ at the center of the Brillouin zone for Be2C and KRbAu2, respectively. The group velocity of Be2C exceeds 10 km/s, while that of KRbAu2 does not exceed 1.1 km/s. This significant difference in group velocity can be attributed to differences in bonding and strength, which is consistent with the inference made from the steepness of their phonon spectra. Fig. 12(c) and (f) describe the differences in phonon lifetimes between the two materials. In the low-frequency range below 10 THz, Be2C exhibits longer phonon lifetimes, with the longest exceeding 10 ps and the shortest being above 1 ps. In contrast, in KRbAu2, the phonon lifetime decreases sharply with increasing frequency, with $\tau_{\lambda}$ approaching 0.1 ps at $\upsilon_{\lambda}=2{\rm THz}$. It is noteworthy that for certain specific low-frequency modes, the phonon lifetime $\tau_{\lambda}$ of KRbAu2 exceeds that of Be2C, which may be due to special scattering suppression effects, e.g., weaker three-phonon scattering [121]. However, overall, Be2C has a longer and more uniform phonon lifetime, which is favorable for long-range heat transport, while the shorter phonon lifetime of KRbAu2 suppresses $\kappa_{\rm L}$.
IV Conclusion
@@
Our work employed KAN to model $\log(\kappa_{\rm L})$, demonstrating that its performance in terms of both accuracy and robustness is fully comparable to that of black-box models. For the interpretability analysis of white-box models, we considered feature interaction effects and dependencies. Sensitivity analysis results based on the KAN and SISSO models revealed that KAN can accurately extract key features of $\kappa_{\rm L}$, highlighting its superior interpretability in this context. We combined KAN with CGCNN to construct a two-stage prediction framework, where CGCNN predicts key features, and the KAN pretraining model maps these features to $\log(\kappa_{\rm L})$; for physical properties like $C_{\rm v}$, which CGCNN struggles to predict accurately, we used DFT calculations to complement the framework. In the case of unlabeled new samples, we applied $k$-means clustering to identify 10 potential thermal conductors/insulators. Among them, the $\kappa_{\rm L}$ values of CN2, CrB4, Be2C, and KRbAu2 were verified by DFT and MD, yielding results highly consistent with deep learning predictions, which sufficiently demonstrate KAN’s excellent extrapolation capability. We visualized the phonon spectra, PDOS, -pCOHP, and other information of Be2C and KRbAu2 using DFT, analyzing the phonon transport mechanisms of both materials from the perspectives of chemical bonding, atomic mass, and structural stability, providing theoretical insights as reference.
V Future perspectives
This work demonstrates the capability of KAN in predicting $\kappa_{\rm L}$, but several key issues remain to be explored for further advancing the research.
Towards an All-Feature Predictor. Among the four parameters describing $\kappa_{\rm L}$, although CGCNN can accurately estimate the vibrational free energy $F$, the static bulk modulus $K_{\rm s}$, and the thermal expansion coefficient $\alpha$ from the crystal structure file, to the best of our knowledge, there is currently no straightforward method to accurately determine the constant-volume heat capacity $C_{\rm v}$ from the crystal structure. Future research should explore more powerful network architectures or models incorporating physical priors to enhance the fitting capability for thermodynamic properties that are difficult to predict, thereby achieving a truly DFT-free, fully machine learning-based high-throughput screening workflow.
Extension to Multi-Temperature Modeling. Due to the limitation of available data, this work focuses on modeling and predicting $\kappa_{\rm L}$ at 300K. In reality, $\kappa_{\rm L}$ is a temperature-dependent variable. Future studies may consider building a comprehensive dataset encompassing temperature variation through DFT or MD calculations or experimental data collection to extend the current framework.
Inverse Design. The framework proposed in this work can be further applied to inverse design tasks: by optimizing the crystal structure, candidate materials with target $\kappa_{\rm L}$ values can be identified. Combined with generative models or evolutionary algorithms, this approach holds promise for the targeted discovery of high or low thermal conductivity materials.
Acknowledgement
We would like to acknowledge the financial support from the Translational Medicine and Interdisciplinary Research Joint Fund of Zhongnan Hospital of Wuhan University (Grant No. ZNJC202235, No. ZNJC202424), the Hubei Provincial Key Technology Foundation of China (No. 2021ACA013), and the National Natural Science Foundation of China (No. 22327901). The numerical calculations in this work have been done on the supercomputing system in the Supercomputing Center of Wuhan University.
Data Availability
@@ Acknowledgement
The original LTC data and the selected features can be downloaded from https://aflowlib.org/ or accessed via the GitHub repository at https://github.com/FlorianTseng/LTC-modeling.
Code Availability
@@
Author Contributions
Yuxuan Zeng: Investigation, Methodology, Writing–original draft, Formal analysis, Data curation, Conceptualization, Visualization. Wei Cao: Writing-original draft, Methodology, Supervision, Resources, Funding acquisition, Project administration. Yijing Zuo: Data curation, Formal analysis, Methodology. Tan Peng: Data curation. Yue Hou: Project administration, Funding acquisition. Ling Miao: Software. Ziyu Wang: Supervision, Resources, Project administration, Funding acquisition, Writing-review & editing. Jing Shi: Resources, Project administration, Funding acquisition.
Competing Interests
The authors declare no competing interests.